OpenStudy (ellecullen):

Homework: Find all intervals where the following function is (strictly) increasing or (strictly) decreasing. f(x) = |x+1|/|x+1| Please help asap. just list the intervals you find, as the answer requires the intervals only.

OpenStudy (anonymous):

so sorry @ElleCullen idk how to do it...:(

OpenStudy (ellecullen):

its ok, I'll keep asking others too :)

OpenStudy (anonymous):

kk!thx for understanding!!

OpenStudy (ellecullen):

I don't think so, as I am doing this math in a super expensive online course. |x+1| = x+1 is the basic absolute value rule.

OpenStudy (ellecullen):

Thanks so much ! :) Elle

OpenStudy (ellecullen):

@Loser66, you get a medal :)

OpenStudy (anonymous):

\[\frac{f(x)}{f(x)}=1\] except for when \(f(x)=0\).